OpenStudy (anonymous):

Solve for L: A = L • W • H L = A times W over H L = W times H over A L = A over quantity W times H L = W over quantity H times A

OpenStudy (anonymous):

it's the second one.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

lets see. i'll plug in #'s, okay? one second.

OpenStudy (anonymous):

A = L * W * H 27 = ? * 3 * 3 3 * 3 ---- = ? 27

OpenStudy (igreen):

A = L * W * H Solve for L: To solve for L you have to divide WH to both sides: That will leave us with A/WH = L So which one's our answer? @mjosedavid

OpenStudy (anonymous):

3 * 3 = 9 27 / 9 = 3

OpenStudy (igreen):

I think your answer is wrong. @silverr

OpenStudy (igreen):

You're going to end up with L = A/WH. That's option C, not option B.

OpenStudy (igreen):

You got the numerator and the denominator mixed up.

OpenStudy (anonymous):

sorry, i wasn't looking close enough on the answer choices? hmm.